OpenStudy (anonymous):

Graphically solve this: |2x-2|=9 I forgot how to do this... help?

OpenStudy (anonymous):

Can I start with taking the absolute lines off and making the -2 a positive? like 2x+2=9 ?

OpenStudy (aum):

Whenever you see |A| = B, then you have to solve two equations without the absolute bars: A = B and A = -B

OpenStudy (anonymous):

okay now it would be 2x+2=9 because its positive ? 2x=7 x=3.5 ?

OpenStudy (aum):

Not 2x+2. 2x - 2.

OpenStudy (anonymous):

are we just assuming its off and then we do calculations and then we put the absolute bars back on?

OpenStudy (anonymous):

at the

OpenStudy (anonymous):

end

OpenStudy (anonymous):

2x-2=9 2x=11 x=5.5?

OpenStudy (aum):

|2x-2|=9 If 2x-2 is positive, the absolute bar does do anything because it is already positive. So for the case 2x-2 > 0: 2x - 2 = 9, 2x = 11, x = 5.5 If 2x-2 is negative, the absolute bar changes it to -(2x-2) and makes it positive. So -(2x-2) = 9 or 2x - 2 = -9; 2x = -7; x = -3.5 So the two solutions are: x = -3.5 and x = 5.5 Try putting each of the value back into the equation and you will see they are true.

OpenStudy (aum):

If 2x-2 is positive, the absolute bar does NOT do anything because it is already positive.

OpenStudy (anonymous):

i thought it changed everything things inside the bars, but i see i guess, so if it was |-2x+2| it would be 2x+2?

OpenStudy (anonymous):

but if it was already |+9999x+9999| it'd stay the same

OpenStudy (anonymous):

wthiuout bars

OpenStudy (anonymous):

9999x+9999

OpenStudy (aum):

No. You have to look at the whole thing and not the sign on individual things. |-2x+2| does not change to 2x+2. You have to consider two cases. If the value of the variable x is such that -2x+2 is positive, then the absolute bars can be dropped because -2x+2 is already positive. The next case is if the value of the variable x is such that -2x+2 is negative, then the absolute bars will change it to -(-2x+2) to make it positive and now the absolute bars can be dropped. The simplest way to remember is: If |A| = B it implies A = B and -A = B or A = -B.

OpenStudy (anonymous):

Oh to find the X's you have to do the 'two case "assuming" method'. I think thats what I was forgetting. Maybe.

OpenStudy (aum):

But this is how I remember it. If you have one equation that has absolute bars you have to solve two equations: |A| = B implies A = B and A = -B |2x-2|=9 implies 2x - 2 = 9 or 2x - 2 = -9 2x = 11 or 2x = -7 x = 5.5 or x = -3.5

OpenStudy (anonymous):

ok I SEe!

OpenStudy (anonymous):

WILL THIS WORK ALL THE TIME?

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

OKAY I did them and i got the same answers as yours, -3.5 and 5.5

OpenStudy (anonymous):

so (-3.5,9) and (5.5,9)

OpenStudy (aum):

And you can verify they are true by plugging them back into the original equation: put x = -3.5 in |2x-2|. |2*(-3.5) - 2| = |-7 - 2| = |-9| = 9 (true) put x = 5.5 in |2x-2|. |2*(5.5) - 2| = |11 - 2| = |9| = 9 (true)

OpenStudy (anonymous):

yes i did that, :D

OpenStudy (aum):

The answer is simply x = -3.5 and x = 5.5 This is an equation with no y-values involved and so don't put them in ordered pair (x,y).

OpenStudy (anonymous):

im supposed to graph it

OpenStudy (anonymous):

i think teacher ment doing the V thing

OpenStudy (anonymous):

cause we did do one of those in class. we are revising for pre calc 30

OpenStudy (anonymous):

we are revising pre calc 20

OpenStudy (aum):

To solve the equation graphically do the following: |2x-2| = 9 |2x - 2| - 9 = 0 Plot y = |2x - 2| - 9. The points where it crosses the x-axis will have y = 0 and those are the solutions.

OpenStudy (aum):



OpenStudy (unklerhaukus):

the two solutions are on a number line (the x-axis)

OpenStudy (anonymous):

sorry laptop died

OpenStudy (anonymous):

if x is 1 y is -9?

OpenStudy (aum):

y = |2x - 2| - 9 put x = 0; y = |0 - 2| - 9 = 2 - 9 = -7. So (0, -7) put x = 1; y = |2 - 2| - 9 = 0 - 9 = -9 So (1, -9)

OpenStudy (anonymous):

it seems odd compared to you graph

OpenStudy (aum):

Mine is a rough sketch not to scale and not done on a graph paper.

OpenStudy (anonymous):

okay. then I get it!

OpenStudy (anonymous):

I have more of these, do you have time? (complicated ones)

OpenStudy (aum):

http://prntscr.com/4jiovf

OpenStudy (anonymous):

o wow! i see

OpenStudy (aum):

Close this question and post it separately. I have to log off now but others may help.

OpenStudy (anonymous):

I was look at wolfram alpha graph it didnt look the same http://www.wolframalpha.com/input/?i=%7C2x-2%7C%3D9

OpenStudy (aum):

http://www.wolframalpha.com/input/?i=y+%3D+ |2x-2|-9

OpenStudy (aum):

Look at the second graph in the above Wolfram link.

OpenStudy (aum):

why greater than 0. You are not solving am inequality here but an equation.

OpenStudy (anonymous):

now I'm wasting both our times. Well I guess good night thanks for helping me!!!!

OpenStudy (aum):

you are welcome. http://www.wolframalpha.com/input/?i=abs%282+x-2%29-9+%3D+0